OpenStudy (anonymous):

giving medal for answer tho.. help peopless ? Maureen cut a 20 cm ribbon into exactly three pieces. The first piece is 3 cm shorter than the second piece. The third piece is 4 cm shorter than the second piece. Find the length of the shortest piece. a. 6 cm c. 8 cm b. 10 cm d. 5 cm

OpenStudy (anonymous):

ok so... set : x= length of 2nd piece in cm x- 3= length of 1st piece in cm x-4= length of 3rd piece in cm then since you know they all came from the 20 cm ribbon you add them together and set them equal to 20 x+ x-3 +x-4 = 20 3x - 7= 20 3x= 27 x= 9 the shortest piece is the third piece which is x- 4 x-4 9-4 = 5 The length of the smallest piece is 5 cm. the other lengths are 9 cm 6 cm

OpenStudy (anonymous):

dang your good!! o,o thanks!

OpenStudy (anonymous):

yw